RC meeting and Run Plan, Friday 29th March

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 3:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Happy Brexit Day No Brexit Day
  • Beam energy measurement was completed yesterday evening - https://logbooks.jlab.org/entry/3673350
  • Unfortunately that's most of the good news out of the way for today
  • Beam very trippy over night (https://logbooks.jlab.org/entry/3673504), MCC resolved this issue early this morning...
  • BUT then there was a new issue, https://logbooks.jlab.org/entry/3673538
  • Beam has been off since roughly 9am this morning due to this issue
  • Beam coming back slowly at around 14:30, once it's stable enough and things look OK we'll begin data taking again
  • HMS Dipole interlocked ~15 minutes ago so currently down again

Current Run Plan

  • We are currently running the dummy target for our first kinematic point at this beam energy, Q2 = 4.4, W = 2.74, low epsilon centre setting
  • Once the dummy target is done (~ 1 hour) we will switch back to LH2 and run until tomorrow day shift (Unless the accelerator fairies are VERY kind to us tonight and we get incredible beam/stats through swing)
  • After this, we switch to the left setting (SHMS moves to 13 degrees) for this kinematic point
  • We'll then run at this point until approximately day shift on Sunday, after this it's onto the next kinematic point (where we'll be for quite a while!)
  • This next setting has a very low SHMS angle (the lowest we can move to remotely), I will give an update on the timing of this as things become clearer over the weekend
  • Whichever shift this ends up on will have to monitor the rotation very vigilantly as it is in progress
  • We will not meet over the weekend unless anything major comes up

Notices in Effect for Shift Crew

  • Please take a screenshot of the spectrometer angles when they are moved to a new position and make a log entry. If working on a kinematic that lasts several days take an image daily.
  • Pay attention to the notices posted in the counting house regarding HMS Q1/Q2.
  • If HMS interlocks when ramping, check the live camera image of the HMS power supplies for smoke/water before clearing the interlock.

Special Requests

  • Alexandre had a short access earlier today during the downtime, he has requested another one early next week during the beam studies/RF recovery down periods.